# Statsformer: Validated Ensemble Learning with LLM-Derived Semantic Priors

### v1

@@ Abstract
We introduce Statsformer, a principled framework for integrating large language model (LLM)-derived knowledge into supervised statistical learning. Existing approaches are limited in adaptability and scope: they either inject LLM guidance as an unvalidated heuristic, which is sensitive to LLM hallucination, or embed semantic information within a single fixed learner. Statsformer overcomes both limitations through a guardrailed ensemble architecture. We embed LLM-derived feature priors within an ensemble of linear and nonlinear learners, adaptively calibrating their influence via cross-validation. This design yields a flexible system with an oracle-style guarantee that it performs no worse than any convex combination of its in-library base learners, up to statistical error. Empirically, informative priors yield consistent performance improvements, while uninformative or misspecified LLM guidance is automatically downweighted, mitigating the impact of hallucinations across a diverse range of prediction tasks.
@@ 1 Introduction
Classical frameworks encode prior knowledge through Bayesian priors, weighted regularization, and hierarchical penalties (Gelman et al., 2013; Tibshirani, 1996b; Zou, 2006; Szafranski et al., 2007). Related paradigms such as learning using privileged information (LUPI) provide foundations for leveraging training-time auxiliary information (Vapnik and Izmailov, 2009; Vapnik and Vashist, 2015), while ensemble aggregation theory provides guarantees for combining learners (Dalalyan and Salmon, 2012; Rigollet and Tsybakov, 2012). In practice, incorporating rich domain knowledge often requires carefully curated, human-specified, feature-level priors, whose construction can be costly and difficult to scale with the number of features. Recent advances in LLMs enable automated elicitation of feature-level priors at scale, but the resulting signals may be noisy or inconsistent (Yao et al., 2024; Huang et al., 2024). A modern framework must therefore adaptively and reliably leverage such priors when informative and discount them when unreliable.
@@ 2.1 Supervised Statistical Learning
often with an additional regularization term $\lambda\,\Omega(\theta)$ (e.g., Tibshirani, 1996a; Zou and Hastie, 2005).
@@
where $\ell$ is a prediction loss and $\Omega$ a regularization term. Depending on how $\{\hat{f}_{j}\}_{j=1}^{J}$ are constructed, ensembles fall into two broad regimes. In homogeneous ensembles, all base learners share the same learning algorithm and model class, and diversity is induced by randomization such as bootstrap resampling, feature subsampling, or random initialization. This setting underlies bagging (Breiman, 1996, 2001), which averages randomized predictors to reduce variance, and boosting (Freund and Schapire, 1996; Friedman, 2001), which constructs predictors sequentially via additive corrections to reduce bias. In heterogeneous ensembles, such as stacking (Wolpert, 1992; LeBlanc and Tibshirani, 1996; van der Laan et al., 2007), the base learners use different algorithms and inductive biases (e.g., linear, tree-based, and neural predictors), and the aggregation rule $F$ is fit using out-of-fold predictions.
@@ 5.1 Implementation Details
Breast Cancer 1,000 1,545 (Stiglic and Kokol, 2010)
@@
Nomao 120 1,000 (Candillier and Lemaire, 2012)
@@
We use four strong base learners that admit feature-level prior injection via the adapters in Section 3: Lasso (Tibshirani, 1996a) (penalty factors), XGBoost (Chen and Guestrin, 2016) (feature sampling weights), random forests (Breiman, 2001) (feature sampling and instance weights), and kernel SVMs (Zhang et al., 2011) (feature scaling). Details of prior injection and per-learner gains are given in Appendix E.
@@
We compare Statsformer against strong tabular learning baselines, including both classical methods and AutoML systems. [^10] Statsformer is not an end-to-end AutoML system and does not perform architecture or pipeline search; AutoML methods serve as widely used reference points. Among AutoML approaches, we include AutoGluon (Erickson et al., 2020), a plug-and-play ensemble of neural networks and tree-based models, and AutoML-Agent (Trirat et al., 2025), a multi-agent LLM framework for full-pipeline AutoML. We also benchmark standard statistical methods, including XGBoost (Chen and Guestrin, 2016), Random Forests (Breiman, 2001), LightGBM (Ke et al., 2017), and Lasso (Tibshirani, 1996a), as well as LLM-Lasso (Zhang et al., 2025), which incorporates LLM-derived feature penalties into Lasso.
@@ 5.2 Results
We ablate the LLM-generated priors to obtain a plain out-of-fold stacking baseline, Statsformer (no-prior), which includes only configurations with $\alpha=\beta=0$. This reduces to a Super Learner-style ensemble that often outperforms individual learners in practice (van der Laan et al., 2007; Polley and van der Laan, 2010), and thus serves as a strong baseline.
@@ Acknowledgment
E.Z., N.S., and F.Z. acknowledge support from the Stanford Graduate Fellowship (SGF) for Sciences and Engineering. N.S. also acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2146755. M.P. acknowledges support in part by National Science Foundation (NSF) CAREER Award under Grant CCF-2236829; in part by the U.S. Army Research Office Early Career Award under Grant W911NF-21-1-0242; in part by the Office of Naval Research under Grant N00014-24-1-2164
@@ A.3 Sufficient conditions for $\Delta_{\mathrm{refit}}\to 0$
Classical algorithmic stability arguments are well suited to Leave-One-Out cross-validation, where removing a single observation produces vanishing perturbations ($O(1/n)$) as $n\to\infty$ [^12] Statsformer applies equally to leave-one-out cross-validation; we use $K$-fold in experiments for computational efficiency, but LOO admits stronger perturbation-based guarantees in certain settings (e.g., for the Lasso (Homrighausen and McDonald, 2013)). . In contrast, for $K$-fold cross-validation with fixed $K$, a constant fraction of the data is removed in each fold. Consequently, the distance $\|\hat{f}_{l}-f_{l}^{(-k)}\|$ does not vanish due to stability alone. Instead, we rely on the consistency of the base learners: if both $\hat{f}_{l}$ and $f_{l}^{(-k)}$ converge to the same fixed population limit $f^{*}_{l}$, their difference must vanish.
@@ A.3.1 Concrete sufficient condition: strongly convex regularized ERM
Let $F(\theta):=\mathbb{E}[\phi(\theta;X,Y)]+\lambda r(\theta)$ and $\hat{F}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\phi(\theta;x_{i},y_{i})+\lambda r(\theta)$. We have that $F$ is $\mu$-strongly convex with unique minimizer $\theta^{*}$. Strong convexity implies $F(\hat{\theta}_{n})-F(\theta^{*})\geq\frac{\mu}{2}\|\hat{\theta}_{n}-\theta^{*}\|_{2}^{2}.$ Moreover, for regularized ERM with $G$-Lipschitz loss (in $\theta$) and $\mu$-strongly convex objective, uniform stability yields the excess risk bound $\mathbb{E}\!\left[F(\hat{\theta}_{n})-F(\theta^{*})\right]\leq\frac{c\,G^{2}}{\mu n}$ for a universal constant $c$ (see (Bousquet and Elisseeff, 2002)). Combining the two displays gives $\mathbb{E}\|\hat{\theta}_{n}-\theta^{*}\|_{2}^{2}\leq\frac{2cG^{2}}{\mu^{2}n}$, hence $\mathbb{E}\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq\frac{\sqrt{2c}\,G}{\mu\sqrt{n}}.$
@@
For estimators such as the Lasso, the tuning parameter may depend on sample size, either through a deterministic schedule $\lambda_{m}$ or data-adaptively via cross-validation. As a result, classical stability arguments based on uniform strong convexity are not the most direct tool for analyzing $K$-fold refitting. In our framework, this is handled by treating the configuration $l$ as including the tuning rule for $\lambda$, and letting $f_{l}^{*}$ denote the corresponding population-limit predictor. Under standard regularity conditions, Lasso predictors tuned by leave-one-out cross-validation are risk-consistent, and the leave-one-out and full-sample fits are asymptotically equivalent (Homrighausen and McDonald, 2013). Consequently, Assumption 2 holds for the Lasso in prediction, and by Lemma 1 we obtain $\Delta_{\mathrm{refit}}\to 0$ (with rates following once $\epsilon_{l}(m)$ is specified).
@@
  • Gradient Boosting: Stage-wise boosting can be made consistent via early stopping, often paired with shrinkage (small step sizes), which implicitly controls the complexity of the additive model in a manner analogous in spirit to the penalty in Proposition 1 (Zhang and Yu, 2005; Bartlett and Traskin, 2007).
@@ Appendix E Base Learner Curation
References (Tibshirani, 1996a) (Chen and Guestrin, 2016) (Breiman, 2001) (Zhang et al., 2011)
@@
We consider gradient-boosted decision trees as implemented in XGBoost (Chen and Guestrin, 2016), a strong nonlinear baseline widely used in tabular learning. While tree-based models do not admit explicit coefficient-level regularization in the same sense as linear models, they naturally allow feature-level prior injection through feature subsampling at decision tree nodes.
@@ F.2 Datasets
The GEMLeR repository (Stiglic and Kokol, 2010) provides a collection of gene expression datasets designed for benchmarking machine learning algorithms on microarray classification problems. Specifically, we use the “OVA Breast Cancer vs. Other” dataset, which has 1545 samples and 10,936 features. Each sample originates from the expO (Expression Project for Oncology) repository, which collects clinically annotated tumor tissue samples processed under standardized conditions using the same microarray platform, with associated clinical outcomes publicly available.
@@
The UCI Nomao dataset (Candillier and Lemaire, 2012) is a benchmark for location de-duplication, where the task is to determine whether a pair of locations corresponds to the same place. It contains 34,465 instances, each representing a pair of locations, described by 120 features comparing attributes such as name, phone number, and geographic information. We use stratified sampling to subsample the dataset to 1,000 instances, preserving the class distribution as described for the Bank Marketing dataset.
@@ F.3 Implementation
The Random Forest and SVM base learners are implemented using scikit-learn (Pedregosa et al., 2011), XGBoost via its Python package, and Lasso via the adelie library for generalized linear models (Yang and Hastie, 2024). For Random Forests and XGBoost, we do not perform per-dataset hyperparameter tuning in order to limit overhead. Since we outperform AutoGluon (Figure 1), which does perform hyperparameter tuning, this choice appears sufficient for the problems considered. Details on each learner, including hyperparameters, are as follows:
@@ G.1.2 Feature-Scaling Adapters: Metric Refinement
Feature-scaling adapters inject priors by rescaling input coordinates before training. For learners whose behavior depends on the input metric (e.g., kernel methods, nearest-neighbor models, and margin-based classifiers), such rescaling can act as a simple form of metric refinement. When geometry is a dominant source of error, a well-aligned scaling can improve conditioning and favorably affect complexity- or margin-based generalization quantities (Bartlett and Mendelson, 2002; Shawe-Taylor and Cristianini, 2004). These effects are typically most relevant in finite-sample regimes and should be interpreted as improved conditioning rather than an expansion of model expressivity.

### v2

@@ Abstract
We introduce Statsformer, a principled framework for integrating large language model (LLM)-derived knowledge into supervised statistical learning. Existing approaches are limited in adaptability and scope: they either inject LLM guidance as an unvalidated heuristic, which is sensitive to LLM hallucination, or embed semantic information within a single fixed learner. Statsformer overcomes both limitations through a guardrailed ensemble architecture. We embed LLM-derived feature priors within an ensemble of linear and nonlinear learners, adaptively calibrating their influence via cross-validation. This design yields a flexible system with an oracle-style guarantee that it performs no worse than any convex combination of its in-library base learners, up to statistical error. Empirically, informative priors yield consistent performance improvements, while uninformative or misspecified LLM guidance is automatically downweighted, mitigating the impact of hallucinations across a diverse range of prediction tasks. An open-source implementation of Statsformer is available at https://github.com/pilancilab/statsformer.
@@ 1 Introduction
Classical frameworks encode prior knowledge through Bayesian priors, weighted regularization, and hierarchical penalties (Gelman et al., 2013; Tibshirani, 1996b; Zou, 2006; Szafranski et al., 2007). Related paradigms such as learning using privileged information (LUPI) provide foundations for leveraging training-time auxiliary information (Vapnik & Izmailov, 2009; Vapnik & Vashist, 2015), while ensemble aggregation theory provides guarantees for combining learners (Dalalyan & Salmon, 2012; Rigollet & Tsybakov, 2012). In practice, incorporating rich domain knowledge often requires carefully curated, human-specified, feature-level priors, whose construction can be costly and difficult to scale with the number of features. Recent advances in LLMs enable automated elicitation of feature-level priors at scale, but the resulting signals may be noisy or inconsistent (Yao et al., 2024; Huang et al., 2024). A modern framework must therefore adaptively and reliably leverage such priors when informative and discount them when unreliable.
@@ 2.1 Supervised Statistical Learning
often with an additional regularization term $\lambda\,\Omega(\theta)$ (e.g., Tibshirani, 1996a; Zou & Hastie, 2005).
@@
where $\ell$ is a prediction loss and $\Omega$ a regularization term. Depending on how $\{\hat{f}_{j}\}_{j=1}^{J}$ are constructed, ensembles fall into two broad regimes. In homogeneous ensembles, all base learners share the same learning algorithm and model class, and diversity is induced by randomization such as bootstrap resampling, feature subsampling, or random initialization. This setting underlies bagging (Breiman, 1996, 2001), which averages randomized predictors to reduce variance, and boosting (Freund & Schapire, 1996; Friedman, 2001), which constructs predictors sequentially via additive corrections to reduce bias. In heterogeneous ensembles, such as stacking (Wolpert, 1992; LeBlanc & Tibshirani, 1996; van der Laan et al., 2007), the base learners use different algorithms and inductive biases (e.g., linear, tree-based, and neural predictors), and the aggregation rule $F$ is fit using out-of-fold predictions.
@@ 5.1 Implementation Details
Breast Cancer 1,000 1,545 (Stiglic & Kokol, 2010)
@@
Nomao 120 1,000 (Candillier & Lemaire, 2012)
@@
We use four strong base learners that admit feature-level prior injection via the adapters in Section 3: Lasso (Tibshirani, 1996a) (penalty factors), XGBoost (Chen & Guestrin, 2016) (feature sampling weights), random forests (Breiman, 2001) (feature sampling and instance weights), and kernel SVMs (Zhang et al., 2011) (feature scaling). Details of prior injection and per-learner gains are given in Appendix E.
@@
We compare Statsformer against strong tabular learning baselines, including both classical methods and AutoML systems. [^10] Statsformer is not an end-to-end AutoML system and does not perform architecture or pipeline search; AutoML methods serve as widely used reference points. Among AutoML approaches, we include AutoGluon (Erickson et al., 2020), a plug-and-play ensemble of neural networks and tree-based models, and AutoML-Agent (Trirat et al., 2025), a multi-agent LLM framework for full-pipeline AutoML. We also benchmark standard statistical methods, including XGBoost (Chen & Guestrin, 2016), Random Forests (Breiman, 2001), LightGBM (Ke et al., 2017), and Lasso (Tibshirani, 1996a), as well as LLM-Lasso (Zhang et al., 2025), which incorporates LLM-derived feature penalties into Lasso.
@@ 5.2 Results
We ablate the LLM-generated priors to obtain a plain out-of-fold stacking baseline, Statsformer (no-prior), which includes only configurations with $\alpha=\beta=0$. This reduces to a Super Learner-style ensemble that often outperforms individual learners in practice (van der Laan et al., 2007; Polley & van der Laan, 2010), and thus serves as a strong baseline.
@@ Acknowledgment
E.Z., N.S., and F.Z. acknowledge support from the Stanford Graduate Fellowship (SGF) for Sciences and Engineering. N.S. also acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2146755. M.P. acknowledges support in part by National Science Foundation (NSF) CAREER Award under Grant CCF-2236829; in part by the U.S. Army Research Office Early Career Award under Grant W911NF-21-1-0242; in part by the Office of Naval Research under Grant N00014-24-1-2164. J.B. gratefully acknowledges support from DoD through the grants Air Force Office of Scientific Research under award number FA9550-20-1-0397 and ONR 1398311, also support from NSF via grants 2229012, 2312204, 2403007 is gratefully acknowledged.
@@ A.3 Sufficient conditions for $\Delta_{\mathrm{refit}}\to 0$
Classical algorithmic stability arguments are well suited to Leave-One-Out cross-validation, where removing a single observation produces vanishing perturbations ($O(1/n)$) as $n\to\infty$ [^12] Statsformer applies equally to leave-one-out cross-validation; we use $K$-fold in experiments for computational efficiency, but LOO admits stronger perturbation-based guarantees in certain settings (e.g., for the Lasso (Homrighausen & McDonald, 2013)). . In contrast, for $K$-fold cross-validation with fixed $K$, a constant fraction of the data is removed in each fold. Consequently, the distance $\|\hat{f}_{l}-f_{l}^{(-k)}\|$ does not vanish due to stability alone. Instead, we rely on the consistency of the base learners: if both $\hat{f}_{l}$ and $f_{l}^{(-k)}$ converge to the same fixed population limit $f^{*}_{l}$, their difference must vanish.
@@ A.3.1 Concrete sufficient condition: strongly convex regularized ERM
Let $F(\theta):=\mathbb{E}[\phi(\theta;X,Y)]+\lambda r(\theta)$ and $\hat{F}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\phi(\theta;x_{i},y_{i})+\lambda r(\theta)$. We have that $F$ is $\mu$-strongly convex with unique minimizer $\theta^{*}$. Strong convexity implies $F(\hat{\theta}_{n})-F(\theta^{*})\geq\frac{\mu}{2}\|\hat{\theta}_{n}-\theta^{*}\|_{2}^{2}.$ Moreover, for regularized ERM with $G$-Lipschitz loss (in $\theta$) and $\mu$-strongly convex objective, uniform stability yields the excess risk bound $\mathbb{E}\!\left[F(\hat{\theta}_{n})-F(\theta^{*})\right]\leq\frac{c\,G^{2}}{\mu n}$ for a universal constant $c$ (see (Bousquet & Elisseeff, 2002)). Combining the two displays gives $\mathbb{E}\|\hat{\theta}_{n}-\theta^{*}\|_{2}^{2}\leq\frac{2cG^{2}}{\mu^{2}n}$, hence $\mathbb{E}\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq\frac{\sqrt{2c}\,G}{\mu\sqrt{n}}.$
@@
For estimators such as the Lasso, the tuning parameter may depend on sample size, either through a deterministic schedule $\lambda_{m}$ or data-adaptively via cross-validation. As a result, classical stability arguments based on uniform strong convexity are not the most direct tool for analyzing $K$-fold refitting. In our framework, this is handled by treating the configuration $l$ as including the tuning rule for $\lambda$, and letting $f_{l}^{*}$ denote the corresponding population-limit predictor. Under standard regularity conditions, Lasso predictors tuned by leave-one-out cross-validation are risk-consistent, and the leave-one-out and full-sample fits are asymptotically equivalent (Homrighausen & McDonald, 2013). Consequently, Assumption 2 holds for the Lasso in prediction, and by Lemma 1 we obtain $\Delta_{\mathrm{refit}}\to 0$ (with rates following once $\epsilon_{l}(m)$ is specified).
@@
  • Gradient Boosting: Stage-wise boosting can be made consistent via early stopping, often paired with shrinkage (small step sizes), which implicitly controls the complexity of the additive model in a manner analogous in spirit to the penalty in Proposition 1 (Zhang & Yu, 2005; Bartlett & Traskin, 2007).
@@ Appendix E Base Learner Curation
References (Tibshirani, 1996a) (Chen & Guestrin, 2016) (Breiman, 2001) (Zhang et al., 2011)
@@
We consider gradient-boosted decision trees as implemented in XGBoost (Chen & Guestrin, 2016), a strong nonlinear baseline widely used in tabular learning. While tree-based models do not admit explicit coefficient-level regularization in the same sense as linear models, they naturally allow feature-level prior injection through feature subsampling at decision tree nodes.
@@ F.2 Datasets
The GEMLeR repository (Stiglic & Kokol, 2010) provides a collection of gene expression datasets designed for benchmarking machine learning algorithms on microarray classification problems. Specifically, we use the “OVA Breast Cancer vs. Other” dataset, which has 1545 samples and 10,936 features. Each sample originates from the expO (Expression Project for Oncology) repository, which collects clinically annotated tumor tissue samples processed under standardized conditions using the same microarray platform, with associated clinical outcomes publicly available.
@@
The UCI Nomao dataset (Candillier & Lemaire, 2012) is a benchmark for location de-duplication, where the task is to determine whether a pair of locations corresponds to the same place. It contains 34,465 instances, each representing a pair of locations, described by 120 features comparing attributes such as name, phone number, and geographic information. We use stratified sampling to subsample the dataset to 1,000 instances, preserving the class distribution as described for the Bank Marketing dataset.
@@ F.3 Implementation
The Random Forest and SVM base learners are implemented using scikit-learn (Pedregosa et al., 2011), XGBoost via its Python package, and Lasso via the adelie library for generalized linear models (Yang & Hastie, 2024). For Random Forests and XGBoost, we do not perform per-dataset hyperparameter tuning in order to limit overhead. Since we outperform AutoGluon (Figure 1), which does perform hyperparameter tuning, this choice appears sufficient for the problems considered. Details on each learner, including hyperparameters, are as follows:
@@ G.1.2 Feature-Scaling Adapters: Metric Refinement
Feature-scaling adapters inject priors by rescaling input coordinates before training. For learners whose behavior depends on the input metric (e.g., kernel methods, nearest-neighbor models, and margin-based classifiers), such rescaling can act as a simple form of metric refinement. When geometry is a dominant source of error, a well-aligned scaling can improve conditioning and favorably affect complexity- or margin-based generalization quantities (Bartlett & Mendelson, 2002; Shawe-Taylor & Cristianini, 2004). These effects are typically most relevant in finite-sample regimes and should be interpreted as improved conditioning rather than an expansion of model expressivity.